A survey of 750 workers were asked how much they use the internet at work. 456 said they used it within limits, and 166 said they did not use the Internet at work.
a. Construct a 95% confidence interval estimate for the proportion of all workers who use it within limits.
b. Construct the same for those who did not use the internet.

Answer by stanbon(75887)

A survey of 750 workers were asked how much they use the internet at work. 456 said they used it within limits, and 166 said they did not use the Internet at work.
a. Construct a 95% confidence interval estimate for the proportion of all workers who use it within limits.
---
p-hat = 456/750 = 0.61
ME = 1.96*sqrt[0.61*0.59/750] = 0.04
--
95% CI: 0.61-0.04 < p < 0.61+0.04
=====================================
Cheers,
Stan H.
b. Construct the same for those who did not use the internet.
Same procedure.